OpenStudy (anonymous):

gcf of 6s^3 t^2 and 16st^4

OpenStudy (anonymous):

2st^2 is the gcf

OpenStudy (anonymous):

Just look for what's common to both....to find the gcf

OpenStudy (anonymous):

Im doing that but some dont come out right. what about 18p^3 q^2,8pq^3 and 20pq^3

OpenStudy (anonymous):

2pq^2

OpenStudy (anonymous):

Can you help me with this equation x(2x-10)(2x+10)=0

OpenStudy (anonymous):

What do you have to do? Find x

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay, so you just simply treat each separately abc = 0.... If abc = 0, and one of the variables equal zero, then they will all equal zero... so we have to find the possible solutions for each x = 0 2x -10 = 0 2x + 10 = 0 Then solve for x for each

OpenStudy (anonymous):

so it would be 0,5.-5.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Would 6e(e+5)=0 would the answer be 0,-5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks